OpenStudy (anonymous):

What is x - 8 < -2 in set notation?

OpenStudy (anonymous):

First, we need to solve for x. x < 6 Converting to set notation this becomes \[x: (-\infty, 6)\] Remember that parenthesis indicate that x cannot equal the value, while square brackets indicate x can equal the value. For example, if we had\[x \le 6\]In set notation, this would be \[x: (-\infty, 6]\]

OpenStudy (anonymous):

lets give each other a medal